if the distance of a falling object is directly proportionae to the square of the time it falls before hitting ground . if the object fell 144 ft in 3 seconds how far will ithave fallen by the end of 4 seconds.how dO I put this as an equation?
I put it together as
if the equation is 144 X 4^2 / 3^2? what would Ido next?

Answer by oscargut(2103)

d=distance
t=time
K=constant of proportion
the distance of a falling object is directly proportionae to the square of the time it falls before hitting ground
then d=Kt^2
the object fell 144 ft in 3 seconds
then 144=K(3)^2
then 144=9K
then K=16
then equation is d=16t^2
so if t =4 d=16(4)^2=16^2=256
Answer: it falls 256 ft after 4 seconds